# Reawakening knowledge: Anticipatory recovery from catastrophic interference via structured training

### v1

@@ Abstract
We explore the training dynamics of neural networks in a structured non-IID setting where documents are presented cyclically in a fixed, repeated sequence. Typically, networks suffer from catastrophic interference when training on a sequence of documents; however, we discover a curious and remarkable property of LLMs finetuned sequentially in this setting: they exhibit anticipatory behavior, recovering from the forgetting on documents before encountering them again. The behavior emerges and becomes more robust as the architecture scales up its number of parameters. Through comprehensive experiments and visualizations, we uncover new insights into training over-parameterized networks in structured environments. Code is available at https://github.com/Agentic-Learning-AI-Lab/anticipatory-recovery-public.
@@ 1 Introduction
(a) Loss on Document 1
(b) Shift-averaged loss
Figure 1: Loss curves for cyclic fine-tuning on the pre-trained Pythia-1b model, with 25 documents. The small black circles indicate points just prior to training on the focal task. These bell-shaped loss curves within each epoch demonstrate the anticipatory recovery phenomenon.
@@
Large language models (LLMs) (Devlin et al., 2019; Brown et al., 2020; Touvron et al., 2023; OpenAI, 2023) have demonstrated remarkable general capabilities in a wide range of natural language tasks. During the training of LLMs, documents are typically uniformly sampled at random. Due to the large scale of the training set—in contrast to many other domains—LLM training typically occurs in an online fashion: each document is used only once for just one update step without further repetition (Hoffmann et al., 2022; Chowdhery et al., 2023; Xue et al., 2024).
Such a training style is in stark contrast with how real world agents like humans acquire new knowledge. In naturalistic settings, the material we’re exposed to is structured in time and often repeats. And given the cost of acquiring information in the real world, people aim to maximize their information gain from each episode. Obtaining new data is often associated with a cost, whether a mental switching cost—as when we go from one lecture to another—or a time cost of waiting for the information. In such scenarios, we hypothesize that real-world agents will spend more compute to best leverage each document than to switch among many.
@@
Toward the goal of investigating more naturalistic training setups, we study a simplistic setting involving structured training of LLMs: documents are presented cyclically in a fixed sequence and are repeated multiple times, just as we humans go through our daily routines. Moreover, to account for the cost of switching among documents, we allow the network take multiple gradient steps for each document.
@@
Typically, networks exhibit catastrophic interference (also known as catastrophic forgetting) (McCloskey and Cohen, 1989) when training on a sequence of tasks: the loss on a given document increases as the training advances to other documents. Surprisingly, we show that in a structured training environment, LLMs exhibit a curious anticipatory recovery behavior: they recover from the forgetting of one document before seeing it again, multiple steps in the sequence prior to the recurrence of the document (see Figure 1). It is analogous to a person anticipating to eat breakfast while taking a morning shower, but leaving the thought aside for the rest of the day. It is remarkable as there is no explicit memory in LLMs that stores sequential knowledge across documents, and there is no systematic overlap of content across documents—the behavior emerges from a random document sequence after repeated exposure to that sequence. Furthermore, only large scale networks reawaken their old knowledge during cyclic training, while smaller ones exhibit no such behavior.
@@
Through extensive experiments, we study how different factors in model architecture and training contribute to the anticipatory recovery phenomenon. We offer insights on the training dynamics in sequentially and cyclically structured input data, and we propose hypotheses for the causes of the behavior. We also show that this phenomenon is not unique to LLMs; some vision models with sufficient width and depth also demonstrate a similar behavior, but LLMs on language modeling tasks exhibit the strongest recovery.
@@ 2 Data and Experiment Setup
In this section we describe the models, datasets, and training setups that we use in the subsequent LLM experiments. Additional details are presented in Appendix A.
@@
We use Pythia (Biderman et al., 2023), a suite of decoder-only autoregressive language models pre-trained on the Pile dataset (Gao et al., 2020; Biderman et al., 2022). The Pythia suite has 8 models of different sizes. For our purposes, we use the five medium-sized models from 160M to 2.8B parameters. We use the fully pre-trained model (at 143K training steps) as well as the untrained initializations to study the effect of pre-training.
@@
We use the CNN/Daily Mail news dataset (Nallapati et al., 2016). The dataset is originally designed for text summarization; we re-purpose it for causal language modeling by discarding the summaries and only using the articles as training data. Importantly, the CNN/Daily Mail dataset is not part of the Pile dataset and hence it is a new domain for the Pythia pre-trained models.
We use the same documents for both training and evaluation. Our goal here is not to determine whether a trained model generalizes to new documents, but rather to study the memory for a particular document as a function of position within the training history.
@@
We randomly sample $T$ documents from the CNN/Daily Mail news dataset. In pre-processing, we truncate each document and take only the first $C$ tokens (we refer to $C$ as “context length” in subsequent text). We then fine-tune our language model on each pre-processed sample for $M$ gradient steps (i.e., using a batch size of 1). We refer to the multiple gradient updates of each document as an “episode”. After each episode we evaluate the loss on all $T$ documents. We repeat the training process for $E$ epochs, where an epoch consists of one episode of each document in a fixed sequence. We use a vanilla gradient descent optimizer with learning rate $0.001$. Unless otherwise stated, the default hyper-parameters in the subsequent experiments are $T=25$, $C=256$, $M=10$, $E=5$. We use the average cross entropy loss (average negative log-likelihood for each token) as our training and evaluation metric. For all experiments we run 3 to 5 trials with different random seeds. The shaded area in the figures denotes standard deviation among trials.
@@ 3 Emergent Anticipatory Recovery
(a) Pre-trained Models
(b) Random Initializations
Figure 2: Effect of model size for (a) pre-trained models and (b) random initializations. In each subfigure, the left shows shift-averaged loss curves for five model sizes and the right shows the recovery score as a function of model size.
In this section, we present our experimental results that reveal the anticipatory recovery phenomenon in the cyclic fine-tuning of large language models. We then demonstrate that anticipatory recovery is an emergent behavior that appears only for models with sufficient capacity.
@@ 3.1 The Anticipatory Recovery Phenomenon
In this first experiment (Figure 1), we have $T=25$ documents, and we do cyclic fine-tuning on the documents for $E=5$ epochs in the same ordering. Both the documents and the ordering are sampled at random beforehand, but kept fixed during the sequential fine-tuning process. We refer to these $T$ documents as $\bm{x}_{1},\cdots,\bm{x}_{T}$. At the start, we fine-tune on document 1 for 10 gradient steps, leading to a significant decrease in the model’s loss on $\bm{x}_{1}$. As we move away from $\bm{x}_{1}$ and fine-tune on other documents, we naturally observe forgetting: the model’s loss on $\bm{x}_{1}$ gradually increases, due to catastrophic interference, until we finish fine-tuning on all other documents and return to $\bm{x}_{1}$. As we iterate through the same document sequence for a second time, we would normally expect the loss on $\bm{x}_{1}$ to still keep increasing after the initial decrease. However, Figure 0(a) shows that the loss on $\bm{x}_{1}$ peaks around $\bm{x}_{15}$ (episode 40) and then starts to decrease. Before we return to $\bm{x}_{1}$, it has recovered more than half of the model’s initial forgetting during the second epoch. We refer to this counter-intuitive decrease in loss as the anticipatory recovery phenomenon. In Figure 0(b), we plot the losses for all the documents and re-align them so that 00 on the x-axis refers to the loss on each document $t$ immediately before training on it for the first time. The figure confirms that the anticipatory recovery phenomenon exists for not only $\bm{x}_{1}$ but all documents.
@@
To quantify the strength of the anticipatory recovery phenomenon, we define the recovery score as the proportion of the initial forgetting during the current epoch that the model recovers before returning to the same document. Mathematically, let the mean (over $t$) of the maximum loss on each document $\bm{x}_{t}$ between the $n^{\text{th}}$ and $(n+1)^{\text{th}}$ time we train on that document be $l_{\text{max}}(n)$, right before the $(n+1)^{\text{th}}$ time we train on it be $l_{\text{before}}(n)$, and right after the $(n+1)^{\text{th}}$ time we train on it be $l_{\text{after}}(n)$. Then we define the recovery score (RS) for epoch $n$ to be [^1] In some cases, a randomly initialized model will produce loss curves that decrease throughout the epoch, because its knowledge is so poor that it enjoys positive generalization among all documents. This yields an infinite or misleadingly large recovery score under Equation 1. We do not include such cases in our experiments so do not bother with more nuanced recovery scores.
$RS(n)=\frac{l_{\text{max}}(n)-l_{\text{before}}(n)}{l_{\text{max}}(n)-l_{\text{after}}(n-1)}.$ (1)
In the following subsections we compute the recovery scores for different model sizes and training hyper-parameters (Figures 2 and 5) to investigate their effects on the anticipatory recovery phenomenon.
@@ 3.2 Anticipatory Recovery is an Emergent Behavior
To study how the model size affects the amount of anticipatory recovery, we repeat this experiment with pre-trained Pythia models (Biderman et al., 2023) of different sizes, specifically 160M, 410M, 1B, and 2.8B. We plot the average loss curves as well as the recovery score for epoch 4 in Figure 1(a). We observe that larger models clearly demonstrate stronger anticipatory recovery. The sharp increase of average recovery score from the 160M model to the 410M model indicates that anticipatory recovery is an emergent behavior.
@@
All our previous experiments are conducted with pre-trained language models. To study whether anticipatory recovery is a result of pre-training, we repeat the experiments on randomly initialized models of different sizes, and plot the loss curves and average recovery scores in Figure 1(b). We follow the model initialization recipe of Biderman et al. (2023). From the loss curves for the 410M and 1B model, especially in the last epoch, we see that the anticipation phenomenon also exists in randomly initialized LLMs. We observe that the anticipation effect is not as strong as in the pre-trained models. The effect of model size still holds: larger models clearly demonstrate stronger recovery.
@@
To further study the effect of model width and depth on the anticipatory recovery phenomenon beyond the model hyperparameters in the Pythia suite, we take a Pythia-1B model and vary the width (size of token embedding) and depth (number of transformer blocks) of the model and plot their average loss curves for cyclic training from random initializations in Figure 4. The original Pythia-1B model has token embedding of size 2048 and 16 transformer blocks. From the plots we observe that the model needs sufficient width (at least 512) and depth (at least 8 transformer blocks) to exhibit noticeable recovery, confirming that it is an emergent behavior contingent on model size.
@@
Refer to caption (a) Size of Token Embedding Refer to caption (b) Number of Model Blocks
Refer to caption (a) Random Masking Refer to caption (b) Random Window Shift
Figure 4: Effect of different levels of data randomization, with (a) random masking with probability up to $0.3$; (b) random shift of the context window up to $128$ tokens.
@@ 3.3 Other Influential Factors
In this section we discuss the effect of other training hyperparameters on the anticipatory recovery phenomenon. In each of the following paragraphs we vary one factor in the training setup and plot the different average loss curves for cyclic fine-tuning of a pre-trained Pythia-1B model. We also include some additional results in Appendix B.
(a) Number of Documents
(b) Number of Gradient Steps
(c) Context Length
(d) Number of Frozen Blocks
@@
We first try to increase the number of documents/tasks ($T$) the model is fine-tuned on in each epoch, with $T\in\{10,25,50,100,200\}$, and plot all the loss curves in Figure 4(a). From the figure we can observe clear recovery for all the curves, suggesting that the model can “memorize” a certain task transition even after training on 200 other tasks.
@@
Figure 4(b) plots training curves with different numbers of gradient steps ($M$) taken on each task ($M\in\{1,5,10,20\}$). More gradient steps in general leads to higher recovery score, although in Appendix B.2 we show that slight anticipation is still observed for $1$ gradient step if we use a larger learning rate. We also demonstrate in the appendix that the anticipation effect is stronger when the same total update is divided among more gradient steps by scaling the learning rate inversely with $M$.
@@
We experiment with different context lengths ($C\in\{128,256,512,1024\}$). Documents are padded to the same length with padding tokens. The loss curves are plotted in Figure 4(c). With the same number of gradient steps, larger context length is correlated with lower recovery score. This suggests that sufficient training on each task is necessary, and for longer input context it takes more gradient descent steps to memorize the task. For example, we show in Section B.6 that we can achieve the recovery score of context length 256 with $M=10$ for context length 1024 with $M=40$.
@@
We experiment with keeping only a subset of transformer blocks in the pre-trained model tunable and freezing the other layers during fine-tuning. Specifically, we freeze the first $B\in\{4,6,8,10,12\}$ transformer blocks (out of 16 total blocks in the Pythia-1B model) and tune only the last $16-B$ blocks. Loss curves are plotted in Figure 4(d). Results show that we need at least 8 tunable transformer blocks to observe a strong anticipation phenomenon in the last epoch. The results are consistent with our observations in section 3.2 and confirm that the model needs sufficient depth to exhibit anticipatory recovery even with a frozen pre-trained deep representation.
@@
In addition to the gradient descent optimizer, we also experiment with the Adam optimizer (Kingma and Ba, 2015). For each document, we reset the optimizer state. Loss curves are plotted in Figure 6. Results show that Adam, which is a stronger optimizer than standard gradient descent, further facilitates anticipatory recovery, at least in part by producing greater initial forgetting. The average recovery score for Adam is $0.689$ while for vanilla gradient descent is $0.211$.
@@
So far we have experimented with the exact same training documents across different epochs. However, in a more realistic sequential learning setting the data points might be slightly different for each repetition (e.g. different descriptions of the same concept, different perspectives of the same object), leading to stochasticity in the optimization process. To explore sequential cyclic training with data randomness, we design the following two training settings: for each gradient step on a document, we experimented with (1) randomly masking a subset (up to 30%) of the tokens in the input, and (2) randomly shifting the “window” of $C$ tokens used for training by up to $C/2$ tokens. The resulting loss curves are plotted in Figure 4. We observe that, while the anticipatory recovery phenomenon is generally weaker when there is more variation in each data point, the recovery still clearly exists.
@@
The experiment results in this subsection suggest that the model’s ability to fit on each task is crucial for the appearance of the anticipatory recovery phenomenon. With a larger number of gradient steps, shorter context length, more learnable layers, and a better optimizer, the model is more capable of fitting to the focal task, and those factors also correlate with larger recovery score. We also confirmed that anticipatory recovery exists for long task sequences and slightly augmented data points within each episode, and again these factors that make learning harder also reduce the anticipatory recovery.
Refer to caption Figure 6: Comparison between Adam and vanilla gradient descent optimizer for randomly initialized Pythia-1B models on cyclic training.
Refer to caption (a) Causal Image Modeling Refer to caption (b) Image Classification
Figure 7: Results for cyclic training on vision tasks: (a) Causal image modeling with Image GPT. (b) image classification with vision transformers and VGG convolutional networks.

### v2

@@ Abstract
We explore the training dynamics of neural networks in a structured non-IID setting where documents are presented cyclically in a fixed, repeated sequence. Typically, networks suffer from catastrophic interference when training on a sequence of documents; however, we discover a curious and remarkable property of LLMs finetuned sequentially in this setting: they exhibit anticipatory behavior, recovering from the forgetting on documents before encountering them again. This behavior occurs even though the documents are never presented in context together. The behavior emerges and becomes more robust as the architecture scales up its number of parameters. Through comprehensive experiments and visualizations, we demonstrate a new mechanism by which over-parametrized neural networks can recover from catastrophic interference and uncover new insights into training over-parameterized networks in cyclically structured environments.
@@ 1 Introduction
Large language models (LLMs) (devlin2018bert, ; brown2020language, ; touvron2023llama, ; openai2023gpt4, ) have demonstrated remarkable general capabilities in a wide range of natural language tasks. During the training of LLMs, documents are typically uniformly sampled at random. Due to the large scale of the training set—in contrast to many other domains—LLM training typically occurs in an online fashion: each document is used only once for just one update step without further repetition (hoffmann2022training, ; chowdhery2023palm, ; xue2023repeat, ).
Such a training style is in stark contrast with how real world agents like humans acquire new knowledge. In naturalistic settings, the material we are exposed to is structured in time and often repeats in predictable, quasi-cyclic patterns (e.g., a person’s everyday morning routine consists of first taking a shower, then eating breakfast, and finally dressing up). Hence it is important to understand how existing deep learning methods and architectures perform in this setting.
@@
Toward the goal of investigating more naturalistic training setups, we study a simplistic setting involving structured training of LLMs: documents are presented cyclically in a fixed sequence and repeated multiple times, just as we humans go through our daily routines. Moreover, to account for the cost of switching among documents (analogous to the mental switching cost between different environments and the waiting cost of obtaining new data), we allow the network to take multiple gradient steps for each document. Compared to standard task-incremental and class-incremental continual learning settings chen2018lifelong which experience each task only once, our cyclic training setting better approximates the quasi-cyclic temporal structure of real-world environments.
Figure 1: (a) Loss curves on document 1 for cyclic and random shuffled fine-tuning on a pre-trained Pythia-1B model. The black circles indicate points just prior to training on the focal document. The inverted-U loss curves within each epoch demonstrate the anticipatory recovery phenomenon. (b) Shift-averaged loss curve for cyclic fine-tuning. (c) Online loss curves for cyclic and random shuffled fine-tuning with prequential evaluation.
Model Size 410M 1B 1.4B 2.8B
Cyclic 1.09 $\pm$ 0.03 1.03 $\pm$ 0.03 1.14 $\pm$ 0.06 1.34 $\pm$ 0.06
Shuffle 1.34 $\pm$ 0.02 1.51 $\pm$ 0.04 1.51 $\pm$ 0.04 1.79 $\pm$ 0.06
Table 1: Average online loss across epochs 2 to 5 for cyclic fine-tuning and random shuffled fine-tuning.
@@
Typically, networks exhibit catastrophic interference (also known as catastrophic forgetting) (mccloskey1989catastrophic, ) when training on a sequence of tasks: the loss on a given document increases as the training advances to other documents. Curiously, we discover that in a structured training environment, LLMs exhibit a remarkable anticipatory recovery behavior: they recover from the forgetting of one document before seeing it again, multiple steps in the sequence prior to the recurrence of the document (see Figure 1(a) and 1(b)). It is analogous to a person anticipating to eat breakfast while taking a morning shower, but leaving the thought aside for the rest of the day. Critically, we never present two documents together in context, so the model cannot directly learn sequential relationships between them. Thus, our finding is surprising as there is no explicit memory in LLMs that stores sequential knowledge across context windows, and there is no systematic overlap of content across documents—the behavior emerges from a random document sequence after repeated exposure to that sequence. Furthermore, we demonstrate that, as a result of anticipatory recovery, training with fixed ordering achieves superior performance than random shuffling in the prequential evaluation (cai2021online, ) setting (see Figure 1(c) and Table 1). For an agent that continuously acts and learns in the real world, the performance on the upcoming task is what matters, and prequential evaluation measures such performance. This result hints at the practical benefits of structured training.
@@
Through extensive experiments, we study how different factors in model architecture and training contribute to the anticipatory recovery phenomenon (Section 3.3). We show that only large-scale networks exhibit this reawakening of knowledge, and smaller ones exhibit no such behavior (Section 3.2). We also show that this phenomenon is not unique to LLMs; some vision models with sufficient width and depth also demonstrate a similar behavior, but LLMs on language modeling tasks exhibit the strongest recovery (Section 3.4). We offer insights on the training dynamics in sequentially and cyclically structured input data and propose hypotheses for the causes of the behavior (Section 4).
@@ 2 Data and Experiment Setup
In this section, we describe the models, datasets, and training setups that we use in the subsequent experiments. Additional details are presented in Appendix A.
@@
For the LLM experiments, we use Pythia (biderman2023pythia, ), a suite of decoder-only autoregressive language models pre-trained on the Pile dataset (gao2020pile, ; biderman2022datasheet, ). We use pre-trained Pythia models ranging from 160M to 2.8B parameters. For the vision experiments, we use pre-trained Image GPT (chen2020generative, ) models for causal image modeling and pre-trained vision transformer (ViT) (dosovitskiy2020image, ) and VGG-19 (simonyan2014very, ) models for image classification.
@@
For the LLM experiments, we use the CNN/Daily Mail news dataset (nallapati2016abstractive, ). We repurpose it for causal language modeling by discarding the summaries and only using the articles. Importantly, the CNN/Daily Mail dataset is not part of the Pile dataset and hence it is a new domain for the Pythia pre-trained models. We use the same documents for both training and evaluation. Our goal here is not to determine whether a trained model generalizes to new documents, but rather to study the memory for a particular document as a function of position within the training history. For the vision experiments, we use images sampled from CIFAR-10 (krizhevsky2009learning, ) and ImageNet (deng2009imagenet, ).
@@
We randomly sample $T$ documents from the dataset. In pre-processing, we truncate each document to the first $C$ tokens (we refer to $C$ as “context length” in subsequent text). We then fine-tune the LLM on each pre-processed sample for $M$ gradient steps (i.e., using a batch size of 1). We refer to the multiple gradient updates of each document as an “episode”. After each episode we evaluate the loss on all $T$ documents. We repeat the training process for $E$ epochs, where an epoch consists of one episode of each document in a fixed sequence. We use a vanilla gradient descent optimizer. Unless otherwise stated, the default hyperparameters in the subsequent experiments are $T=25$, $C=256$, $M=10$, $E=5$. We use the average cross entropy loss (average negative log-likelihood for each token) as our training and evaluation metric.
@@ 3 Emergent Anticipatory Recovery
Figure 2: Effect of model size for (a) pre-trained models and (b) random initializations. In each subfigure, the left shows shift-averaged loss curves and the right shows the recovery score as a function of model size.
In this section, we present our experimental results that reveal the anticipatory recovery phenomenon in cyclic fine-tuning of large language models. We then demonstrate that anticipatory recovery is an emergent behavior that appears only for models with sufficient capacity.
@@ 3.1 The Anticipatory Recovery Phenomenon
In this first experiment, we have $T=100$ documents, and we do cyclic fine-tuning of a pre-trained Pythia-1B model (biderman2023pythia, ) on the documents for $E=5$ epochs in the same ordering. Both the documents and the ordering are sampled at random beforehand, but kept fixed during the sequential fine-tuning process. We refer to these $T$ documents as $\bm{x}_{1},\cdots,\bm{x}_{T}$. At the start, we fine-tune on $\bm{x}_{1}$ for $M=10$ gradient steps, leading to a significant decrease in the model’s loss on $\bm{x}_{1}$. As we move away from $\bm{x}_{1}$ and fine-tune on other documents, we naturally observe catastrophic interference: the model’s loss on $\bm{x}_{1}$ gradually increases until we finish fine-tuning on all other documents and return to $\bm{x}_{1}$. As we iterate through the same document sequence for a second time, we would normally expect the loss on $\bm{x}_{1}$ to increase monotonically after the initial decrease. However, Figure 1(a) shows that the loss on $\bm{x}_{1}$ peaks around $\bm{x}_{60}$ and then starts to decrease. Before we return to $\bm{x}_{1}$, the model has recovered more than half of its initial forgetting during the second epoch. We refer to this counterintuitive decrease in loss as the anticipatory recovery phenomenon. In Figure 1(b), we plot the losses for all the documents and re-align them so that 00 on the x-axis refers to the loss on each document $t$ immediately before training on it for the first time. The figure confirms that the anticipatory recovery phenomenon exists for not only $\bm{x}_{1}$ but all documents. On the other hand, when we randomly shuffle the document order within each epoch (except $\bm{x}_{1}$ is always the first document), we do not observe such anticipatory recovery behavior, and the loss on $\bm{x}_{1}$ keeps increasing before we return to it every time.
@@
To quantify the strength of the anticipatory recovery phenomenon, we define the recovery score as the proportion of the initial forgetting during the current epoch that the model recovers before returning to the same document. Mathematically, let the mean (over $t$) of the maximum loss on each document $\bm{x}_{t}$ between the $n^{\text{th}}$ and $(n+1)^{\text{th}}$ time we train on that document be $l_{\text{max}}(n)$, right before the $(n+1)^{\text{th}}$ time we train on it be $l_{\text{before}}(n)$, and right after the $(n+1)^{\text{th}}$ time we train on it be $l_{\text{after}}(n)$. Then we define the recovery score (RS) for epoch $n$ to be [^1] In some cases, a randomly initialized model will produce loss curves that decrease throughout the epoch, because its knowledge is so poor that it enjoys positive generalization among all documents. This yields a misleadingly large recovery score under this definition. We do not include such cases in our experiments so do not bother with more nuanced recovery scores. $RS(n)=\frac{l_{\text{max}}(n)-l_{\text{before}}(n)}{l_{\text{max}}(n)-l_{\text{after}}(n-1)}.$ In the following subsections we compute the recovery scores for different model sizes and training hyperparameters (Figures 2 and 5) to investigate their effects on the anticipatory recovery phenomenon.
@@ 3.2 Anticipatory Recovery is an Emergent Behavior
To study how the model size affects the amount of anticipatory recovery, we repeat this experiment with pre-trained Pythia models (biderman2023pythia, ) of sizes 160M, 410M, 1.4B, and 2.8B. We plot the average loss curves as well as the recovery score for epoch 4 in Figure 2(a) . We observe that larger models clearly demonstrate stronger anticipatory recovery. The sharp increase of average recovery score from the 160M model to the 410M model indicates that anticipatory recovery is an emergent behavior.
@@
To study whether anticipatory recovery is a result of pre-training, we repeat the experiments on randomly initialized models of different sizes, and plot the loss curves and average recovery scores in Figure 2(b). We follow the model initialization recipe of biderman2023pythia . From the loss curves for the 410M and 1B models, especially in the last epoch, we see that the anticipation phenomenon also exists in randomly initialized LLMs. We observe that the anticipation effect is not as strong as in the pre-trained models. The effect of model size still holds: larger models clearly demonstrate stronger recovery.
@@
To further study the effect of model width and depth on the anticipatory recovery phenomenon beyond the model hyperparameters in the Pythia suite, we take a Pythia-1B model and vary the width (size of token embedding) and depth (number of transformer blocks) of the model and plot the average loss curves for cyclic training from random initializations in Figure 4. The original Pythia-1B model has token embedding of size 2048 and 16 transformer blocks. We observe that the model needs sufficient width (at least 512) and depth (at least 8 transformer blocks) to exhibit noticeable recovery, confirming that it is an emergent behavior contingent on model size.
@@
Figure 4: Effect of data randomization strength. (a) Random masking with probability up to $0.3$; (b) Random shift of context window up to $128$ tokens.
@@ 3.3 Other Influential Factors
In this section we discuss the effect of other training hyperparameters on the anticipatory recovery phenomenon. We also include additional experiment details in Appendix A.1 and additional results in Appendix B.
@@
Figure 5(a) plots the loss curves for different number of documents ($T\in~{}\{10,25,50,100,200\}$). We observe clear recovery for all the curves, suggesting that the model can “memorize” a certain task transition even after training on 200 other tasks.
@@
Figure 5(b) plots training curves with different numbers of gradient steps taken on each document ($M\in\{1,5,10,20\}$). More gradient steps in general leads to a higher recovery score, although in Appendix B.2 and B.8 we show that slight anticipation is still observed for $1$ gradient step if we use a larger learning rate, and that the anticipation effect is stronger when the same total update is divided among more gradient steps by scaling the learning rate inversely with $M$.
@@
Figure 5(c) plots the loss curves for different context lengths ($C\in\{128,256,512,1024\}$). Documents are padded to the same length with padding tokens if they are shorter than the specified context length. With the same number of gradient steps, larger context length is correlated with lower recovery score. This suggests that sufficient training on each task is necessary, and for longer input context it takes more gradient descent steps to memorize the task.
@@
We experimented with freezing the first $B\in\{4,6,8,10,12\}$ transformer blocks in the pre-trained Pythia-1B model and tune only the last $16-B$ blocks. Loss curves are plotted in Figure 5(d) . More frozen transformer blocks is correlated with lower recovery score. This observation is consistent with section 3.2 and confirm that the model needs sufficient depth to exhibit anticipatory recovery even with a frozen pre-trained deep representation.
@@
In addition to the gradient descent optimizer, we experimented with the Adam optimizer (kingma2014adam, ). Loss curves are plotted in Figure 6. We reset the optimizer state for each document. Results show that Adam, which is a stronger optimizer, further facilitates anticipatory recovery for both randomly initialized and pre-trained models.
@@
In realistic sequential learning setting the data points might be slightly different for each repetition (e.g. different descriptions of the same concept, different perspectives of the same object), leading to stochasticity in the optimization process. To explore sequential cyclic training with data randomness, we design the following two training settings: (1) we randomly mask a subset of the tokens in the input, and (2) we randomly shift the “window” of $C$ tokens used for training. The resulting loss curves are plotted in Figure 4. We observe that, while anticipatory recovery is generally weaker when there is more variation in each data point, the recovery still clearly exists.
@@
The experiment results in this subsection suggest that the model’s ability to fit on each task is crucial for the strength of anticipatory recovery. With a larger number of gradient steps, shorter context length, more learnable layers, and a better optimizer, the model is more capable of fitting to the focal task, and those factors also correlate with larger recovery score. We also confirmed that anticipatory recovery exists for long task sequences and slightly augmented data points within each episode, and again these factors that make learning harder also reduce anticipatory recovery.
Figure 6: Comparison between Adam and vanilla gradient descent on (a) randomly initialized and (b) pre-trained Pythia-1B models with cyclic training.
Figure 7: Results for cyclic training on (a) Causal image modeling with Image GPT. (b) image classification with vision transformers and VGG networks.